OpenStudy (anonymous):

Kelly is 5 ft 3 in. tall. At 11 A.M., her shadow is 3 ft 8 in. tall. The shadow of a nearby flagpole is 24 ft 6 in. long. How high is the flagpole? (Round your answer to one decimal place.)

OpenStudy (daggiemib):



OpenStudy (danjs):

similar values

OpenStudy (anonymous):

34.3?

OpenStudy (daggiemib):

cross multiply to solve for x?

OpenStudy (danjs):

the sun will be beaming down at the same angle for both, so you can show similarity for the two

OpenStudy (anonymous):

i got 34.4 when i cross mulitplied

OpenStudy (anonymous):

24.6 X5.3/ 3.8

OpenStudy (danjs):

the corresponding sides for both triangles will be in the same proportion or ratio

OpenStudy (anonymous):

this is hard... its due in like 3 min

OpenStudy (danjs):

it is odd to do math in ft and inches in a single value, i would change it ot one or the other first

OpenStudy (anonymous):

can someone help....

OpenStudy (anonymous):

i need the answer

OpenStudy (danjs):

3'8'' = 3 + (8/12) ft, for example,

OpenStudy (anonymous):

i got .6666 repeating

OpenStudy (danjs):

change the 4 values to ft or in first then put the corresponding sides in ratios that are equal 3'8 / 24 6 = 5 ' 3 / x

OpenStudy (danjs):

3 foot 8 to 24 ft 6 is the same as 5ft 3 to x

OpenStudy (anonymous):

okay so whats x?

OpenStudy (anonymous):

my answer has to be due in 1 min... this is timed

OpenStudy (danjs):

should be near a bit over 30, i didnt calculate sorry

OpenStudy (anonymous):

its saying 34.3 but my homework says its wrong...

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

i just typed it in a calculator, it says 3087/88 which is approx 35.08 ft

OpenStudy (anonymous):

okay thanks so much! thats correct

OpenStudy (danjs):

i put each value as ft 3'8" = 3 + 8/12 ft

OpenStudy (danjs):

yeah, similarity properties for the two triangles, if you care or not what to remember